OpenStudy (anonymous):

Add 4 7/10 and 3 1/10.

OpenStudy (anonymous):

it would be easier to turn these into improper fractions first so 4 7/10= 47/10 and 3 1/10=31/10 then just add across the numerators 47+31=78---->78/10---> simplify and get 39/5